# Optimal approximation of infinite-dimensional holomorphic functions II: recovery from i.i.d. pointwise samples

### v2

@@ 6.3 Reconstruction map: unknown anisotropy
In this section we consider the case where $\bm{b}$ is unknown. Here we need a different approach to the proof of Theorem 4.5, where the construction of $\mathcal{T}$ relied heavily on a fixed $\bm{b}$. Instead, we now build the reconstruction map $\mathcal{T}$ based on the solution of a compressed sensing problem.
@@
$k=\frac{m}{cL},\quad\text{where }L=L(m,\epsilon)=\log^{4+\zeta}(m)+\log(\epsilon^{-1})$ (6.14)
@@
$n=\left\lceil{k^{\log^{\zeta/2}(m)}}\right\rceil,\quad\lambda=\frac{3}{7\sqrt{k}}$ (6.15)
@@
$\begin{split}&\mathcal{T}(\{(\bm{y}_{i},f_{i})\}^{m}_{i=1})=\sum_{\bm{\nu}\in\Lambda}\hat{c}_{\bm{\nu}}\Psi_{\bm{\nu}},\\ &\text{where }\hat{\bm{c}}_{\Lambda}=(\hat{c}_{\bm{\nu}})_{\bm{\nu}\in\Lambda}\in{\underset{\bm{z}\in\mathcal{V}^{N}}{\operatorname{argmin}}}~{}{\Big{\{}}\lambda{\|\bm{z}\|}_{1,\bm{u};\mathcal{V}}+{\|\bm{A}\bm{z}-{m}^{-1/2}\bm{f}\|}_{2;\mathcal{V}}{\Big{\}}},\end{split}$ (6.17)
@@
The proof involves modifying the proof of [3, Thm. 4.2]. A significant difference between these results is that the latter provides a nonuniform guarantee that holds for a fixed function $f$. Moreover, [3, Thm. 4.2] pertains to Chebyshev and Legendre polynomials only. In contrast, we now establish a uniform error bound with high probability that holds for arbitrary Jacobi polynomials. See Remark 6.4 for some further discussion.
@@
Step 1: Establishing the rNSP. Observe that the matrix $\bm{A}$ in (6.3) extends to a bounded linear operator from $\mathcal{V}^{N}$ to $\mathcal{V}^{m}$ in the obvious way. We first show that $\bm{A}$ has the weighted rNSP of order $(k,\bm{u})$ with $\rho=1/2$ and $\gamma=3/2$, where $k>0$ is as in (6.14). By [3, Lem. 6.3], it suffices to show that $\bm{A}\in\mathbb{R}^{m\times N}$ has the weighted RIP over $\mathbb{R}$ of order $(2k,\bm{u})$ with constant $\delta=1/(4\sqrt{2}+1)$. To show this, we apply [3, Lem. 6.4]. Notice that this lemma is stated for Chebyshev and Legendre polynomials only, but it in fact applies to any orthonormal system of functions. This lemma asserts that the desired weighted RIP holds with probability at least $1-\epsilon$, provided
@@
for some universal constant $c_{0}$, where $n$ is as in (6.15). Recall from (6.14) that $k\leq m$ since $m\geq 3$ (which implies that $L\geq 1$) and $c\geq 1$ by assumption. This implies that $n\leq 2m^{\log^{\zeta/2}(m)}$. Hence, with $\delta=1/(4\sqrt{2}+1)$, the right-hand side of the above expression satisfies
$c_{0}\cdot\delta^{-2}\cdot k\cdot\left(\log^{2}(k/\delta)\cdot\log^{2}(\mathrm{e}n)+\log(2/\epsilon)\right)\lesssim k\cdot\left(\log^{4+\zeta}(m)+\log(1/\epsilon)\right)=k\cdot L.$
Therefore, (6.18) holds for sufficiently-large choice of the universal constant $c$ in (6.14).
@@
where $f_{\Lambda}$ is as in (5.3) $\bm{c}_{\Lambda}=(c_{\bm{\nu}})_{\bm{\nu}\in\Lambda}\in\mathcal{V}^{N}$ and $\hat{\bm{c}}_{\Lambda}$ is as in (6.17). Observe that
@@
Recall that, with probability at least $1-\epsilon$, $\bm{A}$ has the weighted rNSP of order $(k,\bm{u})$ with $\rho=1/2$ and $\gamma=3/2$. We now apply [4, Lem. 7.4] with $\bm{w}=\bm{u}$, the given values of $\rho$ and $\gamma$, $\bm{x}=\bm{c}_{\Lambda}$, $\widetilde{\bm{x}}=\hat{\bm{c}}_{\Lambda}$, $\bm{b}=m^{-1/2}(f(\bm{y}_{i}))^{m}_{i=1}$ and $\bm{e}=-\widetilde{\bm{e}}$. Since $\hat{\bm{c}}_{\Lambda}$ is a minimizer of (6.17) and $\lambda$ is defined in (6.15) so that
@@
where $C(\bm{b},p,\tau)=C_{1}(\bm{b},p,\tau)+C_{2}(\bm{b},p,\tau)$, with $C_{i}(\bm{b},p,\tau)$, $i=1,2$, as above. Now suppose that $k\geq 1$. Recall the definition of $n$ in (6.15). Given $10<p<$ 1, there exists an $m_{0}(p,\zeta)$ such that $\log^{\zeta/2}(m)\geq\frac{1/p-1/2}{1/p-1}$ for all $m\geq m_{0}(p,\zeta)$. Therefore
${n\geq k^{\frac{1/p-1/2}{1/p-1}},\quad\forall m\geq m_{0}(p,\zeta).}$
@@
${\|f-\hat{f}\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,{\tau})\cdot\begin{cases}\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}&m\geq m_{0}(p,\zeta)\\ 2&3\leq m<m_{0}(p,\zeta)\end{cases}\leq C(\bm{b},p,\zeta,{\tau})\cdot\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}.$
@@
${\|f-\hat{f}\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,\zeta,{\tau})\cdot\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}.$
@@
    Much as in Remark 6.2, we have fact shown that the constant $C(\bm{b},p,{\zeta,\tau})$ in Theorem 4.6 has the form
    $C(\bm{b},p,{\zeta},{\tau})={C_{0}(p,\zeta)}\cdot\left(C_{1}(\bm{b},p,{\tau})+C_{2}(\bm{b},p,{\tau})\right),$
    where $C_{0}(p,\zeta)$ depends on $p$ and $\zeta$ only, $C_{1}(\bm{b},p,{\tau})$ is the constant of Theorem 5.1, part (a) with $q=1$ and $C_{2}(\bm{b},p,{\tau})$ is the constant of Theorem 5.2, part (a) with $q=1$. We will use this in the second part of the proof of Theorem 4.4 below.
@@
    In order to obtain a uniform guarantee, in the above proof we bound the term ${\|\widetilde{\bm{e}}\|}_{2;\mathcal{V}}$ in (6.20) using the crude, but deterministic bound ${\|\widetilde{\bm{e}}\|}_{2;\mathcal{V}}\leq{\|f-f_{\Lambda}\|}_{L^{\infty}(\mathcal{U};\mathcal{V})}$. In contrast, the proof of [3, Thm. 4.2] uses the fact that ${\|\widetilde{\bm{e}}\|}^{2}_{2;\mathcal{V}}$ is, much as in Remark 6.2, a sum of independent random variables with mean ${\|f-f_{\Lambda}\|}^{2}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}$ in order to provide a sharper probabilistic bound using Bernstein’s inequality. The advantage of this approach is it allows for the smaller choice $n\approx m/(c_{0}L)$ in Step 2. Conversely, in the above proof, to compensate for the cruder bound (6.20) we need to choose $n$ larger, as in (6.15). This has a practical downside, since the computational cost of implementing the reconstruction map $\mathcal{T}$ scales linearly in $N=|\Lambda|=|\Lambda^{\mathsf{HCI}}_{n}|$ (recall that the optimization problem in (6.17) is over $\mathcal{V}^{N}$). It is currently unclear whether uniform bounds with the same rate of convergence in Theorem 4.6 could be obtained with a smaller choice of $n$.
  • Remark 6.5
    It is evident from the above proof that the term $\log^{4+\zeta}(m)$ in Theorem 4.4 could be replaced by $\log^{4}(m)g(m)$ for any increasing function $g:\mathbb{N}\rightarrow(0,\infty)$ with $g(m)\rightarrow+\infty$ as $m\rightarrow+\infty$. Indeed, we simply replace the exponent $\log^{\zeta/2}(m)$ in (6.15) by $\sqrt{g(m)}$ and argue in the same way.
7 Proofs of Theorems 4.3 and 4.4
@@
We now prove the second part of the theorem. Let $\bm{b}\in\ell^{p}(\mathbb{N})$, $\bm{b}\geq\bm{0}$ with ${\|\bm{b}\|}_{p}\leq 1$ and $t\in(p,1)$. Let $C(\bm{b}_{\pi},t,\tau)$ be the constant of Theorem 4.5 with $p$ replaced by $t$ (since $\bm{b}\in\ell^{t}(\mathbb{N})$ as $t>p$), and recall from from Remark 6.2 that the constant $C(\bm{b}_{\pi},t,\tau)=C_{0}(p,t)\cdot C_{1}(\bm{b}_{\pi},t,\tau)$, where $C_{0}(p,t)$ depends on $t$ only and $C_{1}(\bm{b}_{\pi},t,\tau)$ is the constant from Theorem 5.1, part (a) with $q=2$, $p$ replaced by $t$ and with parameters $\bm{\alpha}_{\pi}$, $\bm{\beta}_{\pi}$ and measure $\varrho_{\bm{\alpha}_{\pi},\bm{\beta}_{\pi}}$. Applying Theorem 5.1, part (b) we deduce that
@@
$E=\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {0<p<1}\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,{\zeta}{,}{\tau})\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/p}\right\}.$
Consider arbitrary $10<p<$ 1, $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$ and $f\in\mathcal{H}(\bm{b})$. If the event $E$ occurs, then we have
${\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,{\zeta},{\tau})\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/p}.$
@@
Having bounded the reconstruction error when $E$ does or does not occur, we now argue as in the proof of Theorem 4.3 by law of total expectation to obtain the final result.
@@
$\widetilde{E}=\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{t}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {t\in(p,1)}\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},t,{\zeta},{\tau})\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/t}\right\}.$
From Remark 6.4, we note that $C(\bm{b},t,{\zeta},{\tau})={C_{0}(t,\zeta)}\cdot(C_{1}(\bm{b},t,{\tau})+C_{2}(\bm{b},t,{\tau}))$, where $C_{1}(\bm{b},t,{\tau})$ is the constant of Theorem 5.1, part (a) with $q=1$ and $C_{2}(\bm{b},t,{\tau})$ is the constant of Theorem 5.2, part (a) with $q=1$. Now for any $t\in(p,1)$ applying Theorem 5.1, part (b) and Theorem 5.2, part (b) with $r=t$, we get that
$\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C(\bm{b},t,{\zeta,\tau})\leq{C_{0}(t,\zeta)}\cdot\left(\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C_{1}(\bm{b},t,{\tau})+\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C_{2}(\bm{b},t,{\tau})\right)\leq\widetilde{C}(p,t,{\zeta},{\tau}).$ (7.7)
@@
$\widetilde{E}$ $\subseteq\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {{\|\bm{b}\|}_{p,\mathsf{M}}\leq 1}\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},t,{\zeta},{\tau})\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/t}\right\}$
$\subseteq\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {{\|\bm{b}\|}_{p,\mathsf{M}}\leq 1}\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq{\widetilde{C}(p,t,\zeta,\tau)}\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/t}\right\}$
=$=\left\{\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq{\widetilde{C}(p,t,\zeta,\tau)}\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/t}\right\}=:E.$
@@
$\theta(p,\mathsf{M})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq{\widetilde{C}(p,t,\zeta,\tau)}\left(\frac{m}{\log^{{4+\zeta}}(m)}\right)^{1/2-1/t}.$
To remove the $\log^{{4+\zeta}}(m)$ factor from the bound we note that $t\in{(p,1)}$ is arbitrary. Then, up to a possible change in the constant ${\widetilde{C}(p,t,\zeta,\tau)}$, the result follows immediately. ∎

### v3

@@ 6.3 Reconstruction map: unknown anisotropy
In this section we consider the case where $\bm{b}$ is unknown. Here we need a different approach to the proof of Theorem 4.5, where the construction of $\mathcal{T}$ relied heavily on a fixed $\bm{b}$. Instead, we now build the reconstruction map $\mathcal{T}$ based on the solution of a compressed sensing problem. Since it presents few additional difficulties, we now describe the reconstruction map that achieves the factor $\log^{4}(m)g(m)$ described in Remark 4.3, where $g$ is any function $g:\mathbb{N}\rightarrow[1,\infty)$ with $g(m)\rightarrow+\infty$ subalgebraically fast as $m\rightarrow+\infty$. Theorems 4.3 and 4.6 then follow as special cases with $g(m)=\log(m)$.
@@
$k=\frac{m}{cL},\quad\text{where }L=L(m,\epsilon)=\log^{4}(m)g(m)+\log(\epsilon^{-1})$ (6.14)
@@
$n=\left\lceil k^{\sqrt{g(m)}}\right\rceil,\quad\lambda=\frac{3}{7\sqrt{k}}$ (6.15)
@@
$\begin{split}&\mathcal{T}(\{(\bm{y}_{i},f_{i})\}^{m}_{i=1})=\sum_{\bm{\nu}\in\Lambda}\hat{c}_{\bm{\nu}}\Psi_{\bm{\nu}},\\ &\text{where }\hat{\bm{c}}_{\Lambda}=(\hat{c}_{\bm{\nu}})_{\bm{\nu}\in\Lambda}\in{\underset{\bm{z}\in\mathcal{V}^{N}}{\operatorname{argmin}}}~{}\Big{\{}\lambda{\|\bm{z}\|}_{1,\bm{u};\mathcal{V}}+{\|\bm{A}\bm{z}-m^{-1/2}\bm{f}\|}_{2;\mathcal{V}}\Big{\}}\end{split}$ (6.17)
@@
The proof involves modifying the proof of [4, Thm. 4.2]. A significant difference between these results is that the latter provides a nonuniform guarantee that holds for a fixed function $f$. Moreover, [4, Thm. 4.2] pertains to Chebyshev and Legendre polynomials only. In contrast, we now establish a uniform error bound with high probability that holds for arbitrary Jacobi polynomials. See Remark 6.4 for some further discussion.
@@
Step 1: Establishing the rNSP. Observe that the matrix $\bm{A}$ in (6.3) extends to a bounded linear operator from $\mathcal{V}^{N}$ to $\mathcal{V}^{m}$ in the obvious way. We first show that $\bm{A}$ has the weighted rNSP of order $(k,\bm{u})$ with $\rho=1/2$ and $\gamma=3/2$, where $k>0$ is as in (6.14). By [4, Lem. 6.3], it suffices to show that $\bm{A}\in\mathbb{R}^{m\times N}$ has the weighted RIP over $\mathbb{R}$ of order $(2k,\bm{u})$ with constant $\delta=1/(4\sqrt{2}+1)$. To show this, we apply [4, Lem. 6.4]. Notice that this lemma is stated for Chebyshev and Legendre polynomials only, but it in fact applies to any orthonormal system of functions. This lemma asserts that the desired weighted RIP holds with probability at least $1-\epsilon$, provided
@@
for some universal constant $c_{0}$, where $n$ is as in (6.15). Recall from (6.14) that $k\leq m$ since $m\geq 3$ (which implies that $L\geq 1$) and $c\geq 1$ by assumption. This implies that $n\leq 2m^{\sqrt{g(m)}}$. Hence, with $\delta=1/(4\sqrt{2}+1)$, the right-hand side of the above expression satisfies
$c_{0}\cdot\delta^{-2}\cdot k\cdot\left(\log^{2}(k/\delta)\cdot\log^{2}(\mathrm{e}n)+\log(2/\epsilon)\right)\lesssim k\cdot\left(\log^{4}(m)g(m)+\log(1/\epsilon)\right)=k\cdot L.$
Therefore, (6.18) holds for a sufficiently-large choice of the universal constant $c$ in (6.14).
@@
where $f_{\Lambda}$ is as in (5.5), $\bm{c}_{\Lambda}=(c_{\bm{\nu}})_{\bm{\nu}\in\Lambda}\in\mathcal{V}^{N}$ and $\hat{\bm{c}}_{\Lambda}$ is as in (6.17). Observe that
@@
Recall that, with probability at least $1-\epsilon$, $\bm{A}$ has the weighted rNSP of order $(k,\bm{u})$ with $\rho=1/2$ and $\gamma=3/2$. We now apply [3, Lem. 7.4] with $\bm{w}=\bm{u}$, the given values of $\rho$ and $\gamma$, $\bm{x}=\bm{c}_{\Lambda}$, $\widetilde{\bm{x}}=\hat{\bm{c}}_{\Lambda}$, $\bm{b}=m^{-1/2}(f(\bm{y}_{i}))^{m}_{i=1}$ and $\bm{e}=-\widetilde{\bm{e}}$. Since $\hat{\bm{c}}_{\Lambda}$ is a minimizer of (6.17) and $\lambda$ is defined in (6.15) so that
@@
where $C(\bm{b},p,\tau)=C_{1}(\bm{b},p,\tau)+C_{2}(\bm{b},p,\tau)$, with $C_{i}(\bm{b},p,\tau)$, $i=1,2$, as above. Now suppose that $k\geq 1$. Recall the definition of $n$ in (6.15). Given $10<p<$ 1, there exists an $m_{0}(p,g)$ such that $\sqrt{g(m)}\geq\frac{1/p-1/2}{1/p-1}$ for all $m\geq m_{0}(p,g)$. Therefore
$n\geq k^{\frac{1/p-1/2}{1/p-1}},\quad\forall m\geq m_{0}(p,g).$
@@
${\|f-\hat{f}\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,\tau)\cdot\begin{cases}\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}&m\geq m_{0}(p,g)\\ 2&3\leq m<m_{0}(p,g)\end{cases}\leq C(\bm{b},p,g,\tau)\cdot\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}.$
@@
${\|f-\hat{f}\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,g,\tau)\cdot\left(\frac{m}{L}\right)^{\frac{1}{2}-\frac{1}{p}}.$
@@
    Much as in Remark 6.2, we have in fact shown that the constant $C(\bm{b},p,g,\tau)$ in Theorem 4.6 has the form
    $C(\bm{b},p,g,\tau)=C_{0}(p,g)\cdot\left(C_{1}(\bm{b},p,\tau)+C_{2}(\bm{b},p,\tau)\right),$
    where $C_{0}(p,g)$ depends on $p$ and $g$ only, $C_{1}(\bm{b},p,\tau)$ is the constant of Theorem 5.2, part (a) with $q=1$ and $C_{2}(\bm{b},p,\tau)$ is the constant of Theorem 5.3, part (a) with $q=1$. We will use this in the second part of the proof of Theorem 4.3 below.
@@
    In order to obtain a uniform guarantee, in the above proof we bound the term ${\|\widetilde{\bm{e}}\|}_{2;\mathcal{V}}$ in (6.20) using the crude, but deterministic bound ${\|\widetilde{\bm{e}}\|}_{2;\mathcal{V}}\leq{\|f-f_{\Lambda}\|}_{L^{\infty}(\mathcal{U};\mathcal{V})}$. In contrast, the proof of [4, Thm. 4.2] uses the fact that ${\|\widetilde{\bm{e}}\|}^{2}_{2;\mathcal{V}}$ is, much as in Remark 6.2, a sum of independent random variables with mean ${\|f-f_{\Lambda}\|}^{2}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}$ in order to provide a sharper probabilistic bound using Bernstein’s inequality. The advantage of this approach is that it allows for the smaller choice $n\approx m/(c_{0}L)$ in Step 2. Conversely, in the above proof, to compensate for the cruder bound (6.20) we need to choose $n$ larger, as in (6.15). This has a practical downside, since the computational cost of implementing the reconstruction map $\mathcal{T}$ scales linearly in $N=|\Lambda|=|\Lambda^{\mathsf{HCI}}_{n}|$ (recall that the optimization problem in (6.17) is over $\mathcal{V}^{N}$). It is currently unclear whether uniform bounds with the same rate of convergence as in Theorem 4.6 could be obtained with a smaller choice of $n$.
7 Proofs of Theorems 4.2 and 4.3
@@
We now prove the second part of the theorem. Let $\bm{b}\in\ell^{p}(\mathbb{N})$, $\bm{b}\geq\bm{0}$ with ${\|\bm{b}\|}_{p}\leq 1$ and $t\in(p,1)$. Let $C(\bm{b}_{\pi},t,\tau)$ be the constant of Theorem 4.5 with $p$ replaced by $t$ (since $\bm{b}\in\ell^{t}(\mathbb{N})$ as $t>p$), and recall from Remark 6.2 that the constant $C(\bm{b}_{\pi},t,\tau)=C_{0}(t)\cdot C_{1}(\bm{b}_{\pi},t,\tau)$, where $C_{0}(t)$ depends on $t$ only and $C_{1}(\bm{b}_{\pi},t,\tau)$ is the constant from Theorem 5.2, part (a) with $q=2$, $p$ replaced by $t$ and with parameters $\bm{\alpha}_{\pi}$, $\bm{\beta}_{\pi}$ and measure $\varrho_{\bm{\alpha}_{\pi},\bm{\beta}_{\pi}}$. Applying Theorem 5.2, part (b) we deduce that
@@
$E=\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ 0<p<1\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/p}\right\}.$
Consider arbitrary $10<p<$ 1, $\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N})$, $\bm{b}\geq\bm{0}$, and $f\in\mathcal{H}(\bm{b})$. If the event $E$ occurs, then we have
${\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},p,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/p}.$
@@
Having bounded the reconstruction error when $E$ does or does not occur, we now argue as in the proof of Theorem 4.2 by the law of total expectation to obtain the final result.
@@
$\widetilde{E}=\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{t}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ t\in(p,1)\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},t,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/t}\right\}.$
From Remark 6.4, we note that $C(\bm{b},t,g,\tau)=C_{0}(t,g)\cdot(C_{1}(\bm{b},t,\tau)+C_{2}(\bm{b},t,\tau))$, where $C_{1}(\bm{b},t,\tau)$ is the constant of Theorem 5.2, part (a) with $q=1$ and $C_{2}(\bm{b},t,\tau)$ is the constant of Theorem 5.3, part (a) with $q=1$. Now for any $t\in(p,1)$ applying Theorem 5.2, part (b) and Theorem 5.3, part (b) with $r=t$, we get that
$\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C(\bm{b},t,g,\tau)\leq C_{0}(t,g)\cdot\left(\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C_{1}(\bm{b},t,\tau)+\sup_{\|\bm{b}\|_{p,\mathsf{M}}\leq 1}C_{2}(\bm{b},t,\tau)\right)\leq\widetilde{C}(p,t,g,\tau).$ (7.7)
@@
$\widetilde{E}$ $\subseteq\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {\|\bm{b}\|}_{p,\mathsf{M}}\leq 1\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq C(\bm{b},t,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/t}\right\}$
$\subseteq\bigcap_{\begin{subarray}{c}\bm{b}\in\ell^{p}_{\mathsf{M}}(\mathbb{N}),\bm{b}\geq 0\\ {\|\bm{b}\|}_{p,\mathsf{M}}\leq 1\end{subarray}}\left\{\sup_{f\in\mathcal{H}(\bm{b})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq\widetilde{C}(p,t,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/t}\right\}$
=$=\left\{\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\leq\widetilde{C}(p,t,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/t}\right\}=:E.$
@@
$\theta(p,\mathsf{M})\leq\mathbb{E}\left(\sup_{f\in\mathcal{H}(p,\mathsf{M})}{\|f-\mathcal{T}(\{(\bm{y}_{i},f(\bm{y}_{i}))\}^{m}_{i=1})\|}_{L^{2}_{\varrho}(\mathcal{U};\mathcal{V})}\right)\leq\widetilde{C}(p,t,g,\tau)\left(\frac{m}{\log^{4}(m)g(m)}\right)^{1/2-1/t}.$
To remove the $\log^{4}(m)g(m)$ factor from the bound we note that $t\in(p,1)$ is arbitrary and $g$ grows subalgebraically fast in $m$. Then, up to a possible change in the constant $\widetilde{C}(p,t,g,\tau)$, the result follows immediately. ∎